A Saaeplet ananed and fanto total ayht. of meth cuno expe be between 40e and 440g/nd be beth Cth Hoa h
Answer #1

Ans :

Let the mass of the sample be 100 g

So mass of C = 69.21 g

mol C = mass / molar mass

= 69.21 / 12.01 = 5.76 mol

mass of H = 7.74 g

mol H = 7.74 / 1.008 = 7.68 mol

mass O = 23.05 g

mol O = 23.05 / 15.999 = 1.44 mol

Dividing the number of mol by a common value , we get the ratio of number of mol as :

= 12 : 16 : 3

So the empirical formula becomes : C12H16O3

empirical formula mass = 12(12.01) + 16(1.008) + 3(15.999)

= 208.25 g/mol

Now since the molecular formula lies in the range of 400 - 440 g/mol , we possible believe that it can be generated by multiplying the empirical mass by 2.

So the molecular formula of the compound will be :

= 2 ( C12H16O3)

= C24H32O6